OpenStudy (anonymous):

With the system of equation given what is x? 3x+5y=-9 2x+y=-34

OpenStudy (witty_roger):

x=23

OpenStudy (witty_roger):

sorry -23

OpenStudy (anonymous):

so if thats true i can substitute 3(-23)+5y=-9 to find y?

OpenStudy (witty_roger):

yess

OpenStudy (witty_roger):

obviously..dats one of the standard method of solving linear equations

OpenStudy (witty_roger):

medal plz

OpenStudy (anonymous):

-69+5y=-9 so add 69 to both sides 5y=60 y=12

OpenStudy (witty_roger):

yeah correct..even i got y=12

OpenStudy (anonymous):

lol when i did that it ddnt work lol thank

OpenStudy (witty_roger):

wats going on with @Data_LG2 ..still replying..lol

OpenStudy (anonymous):

lol idk

OpenStudy (anonymous):

Hi, If in case you wanted to learn how it is done, read my comment until the end :) In solving system of linear equations, there are two methods that are usually used. One is the substitution method, in which you have to isolate of your equations in terms of a variable and substitute this variable to your second equation. 3x+5y=-9 --> Equation 1 2x+y=-34 --> Equation 2 Let's arrange equation two in terms of y by subtracting both sides by "2x", it will be \(\sf \color{red}{y= -34 -2x}\). Then, use this equation to solve for x by substituting this y variable from Equation 1: \(\sf 3x+5 \color{red}{y}=-9\\3x+5 \color{red}{( -34 -2x)}=-9\) From this you can easily solve for x. After you solve for x, sub it back to a ny of the equations to solve for y. Next method is the elimination method. In this method you have to manipulate the equation in order to eliminate one variable. Let's take your problem again as an example. 3x+5y=-9 --> Equation 1 2x+y=-34 --> Equation 2 You can multiply all terms in equation 2 by "-5" and add it to equation 1 in order to eliminate variable "y". -5(2x+y=-34) 10x-5y=170 then add it to equation 1. 3x+5y=-9 10x-5y=170 ----------------- 13x = 161 Solve for x and like before, sub the value of x to any of the equations to solve for y :)

OpenStudy (anonymous):

thank you all for your answers :) i wish could give best answer to everyone